OpenStudy (anonymous):

What is the volume of a sphere with a radius of 3 mm? Use 3.14 for π. Round to the nearest hundredth, if needed.

OpenStudy (johnweldon1993):

The formula of Volume for a sphere is (4/3) pi * r^3 So here..3.14 = pi and r = 3mm so Volume = (4/3)(3.14)(3^3 mm) Volume = 339.12/3 mm^3 Volume = 113.04 mm^3

OpenStudy (anonymous):

Thank you:)

OpenStudy (johnweldon1993):

Anytime :)